OpenStudy (summerncaela):

Write an equation in slope-intercept form for the line that contains each pair of points.

OpenStudy (summerncaela):

Here is a pair of points, (3,4),(-1,-2)

OpenStudy (summerncaela):

please help

OpenStudy (ageta):

so are you trying to find the slope

OpenStudy (summerncaela):

Well, I did y2-y1/x2-x1 but thats not the answer

OpenStudy (summerncaela):

the answer in the back of the book says its y=3/2x-1/2 but idk how they got it

OpenStudy (ageta):

they got 3/2x by using the y2-y1/x2-x1

OpenStudy (summerncaela):

how?

OpenStudy (ageta):

just plug them in what is y2

OpenStudy (summerncaela):

4

OpenStudy (ageta):

y2-y1

OpenStudy (summerncaela):

i did 4--2/3--1 =6/4

OpenStudy (ageta):

simplify 6/4 =3/2

OpenStudy (summerncaela):

;l

OpenStudy (summerncaela):

where did the 1/2 come from

OpenStudy (ageta):

use the slope formula now

OpenStudy (ageta):

y=mx+b

OpenStudy (summerncaela):

do i plug in 3/2?

OpenStudy (ageta):

yes

OpenStudy (ageta):

3/2 goes on m which is the slope

OpenStudy (summerncaela):

am a little lost on how to do that

OpenStudy (summerncaela):

oh

OpenStudy (summerncaela):

y=3/2x+b

OpenStudy (ageta):

yes now you need to plug in one cordenate in that formula the x an y

OpenStudy (summerncaela):

i dont understand

OpenStudy (ageta):

from the start you had 2 coordinates right

OpenStudy (summerncaela):

yes

OpenStudy (ageta):

yes one of the then plug them in

OpenStudy (ageta):

x on x and y on y

OpenStudy (ageta):

4=3/2*3+b

OpenStudy (ageta):

because i used (3,4)

OpenStudy (ageta):

now solve for b which is the y-intercept

OpenStudy (summerncaela):

4= 4.5+b then i solve for b?

OpenStudy (ageta):

first don't mix decimals and fractions we want to get fraction so we leave is as fraction

OpenStudy (ageta):

which is 3/2 * 3/1 = 9/2

OpenStudy (ageta):

then substract 9/2 from 4/1 right

OpenStudy (ageta):

b= -1/2

OpenStudy (summerncaela):

wait where did 4/1 come from

OpenStudy (ageta):

4/1 is another way of saying 4 it was on the outside remember

OpenStudy (summerncaela):

it was 4=9/2x+b

OpenStudy (ageta):

is the the Y on (3,4) Y=mx+b

OpenStudy (summerncaela):

i feel so stupid its just not clicking in my brain

OpenStudy (ageta):

4=9/2+b

OpenStudy (summerncaela):

so then you subtract 9/2

OpenStudy (ageta):

we already used the x to multiply 3/2*3 =9/2

OpenStudy (summerncaela):

from both sides?

OpenStudy (ageta):

yes

OpenStudy (ageta):

we want to make b by its self so usbstract 9/2

OpenStudy (summerncaela):

so then its y=9/2x -1/2

OpenStudy (ageta):

no

OpenStudy (summerncaela):

3/2 myy

OpenStudy (ageta):

3/2 is the slope

OpenStudy (summerncaela):

3/2 yes my bad

OpenStudy (ageta):

y=3/2-1/2

OpenStudy (summerncaela):

thank you sso much

OpenStudy (ageta):

i mean y=3/2x-1/2

OpenStudy (ageta):

your welcome

OpenStudy (summerncaela):

you're a life saver tysm

OpenStudy (ageta):

here is helpfull video if you don't get it http://www.virtualnerd.com/algebra-1/linear-equation-analysis/y-intercept-from-point-slope-form.php

OpenStudy (ageta):

it explains it better

OpenStudy (ageta):

bye

OpenStudy (summerncaela):

ok thank u